# Potato Potahto in the FAO-GAEZ Productivity Measures?

### v3

@@ Introduction
$\Delta y_{jt}=\alpha_{t}+\beta\cdot\Delta A_{jt}^{soy}+\sum_{l=1}^{L}\alpha_{l}\cdot W_{l,j}+\epsilon_{jt},$ (9)
where $y_{jt}$ is the employment share of the manufacturing sector in municipality $j$ and year $t$, and $\Delta A_{jt}^{soy}$ is the soybean productivity shock. Bustos et al. (2016) exploits an arguably exogenous shock in soybean productivity: the introduction of genetically modified seeds in 2003, i.e., they use this increase in productivity as an exogenous change in soybean productivity.
@@
We begin by discussing the differences in soybean productivity in versions 3 and 4. We are able to reproduce the numbers of productivity using Bustos et al. (2016)’s replication package. The correlation between the change in productivity variable in the replication package and the one generated by us is of 0.99.131313There are many reasons for differences in the computed productivity. In the replication package, there is no mention of how the variables were transformed from raster and vector formats to the tabular format the authors use. We use the geometry of the minimum comparable areas that is present in the replication package in order to compute the average productivity within each municipality. The replication package does not have the raw soybean productivity data, but we know of no different vintages of the FAO-GAEZ version 3 and, therefore, we have no reason to believe that the data is different. The differences then are likely to occur from different ways of assigning pixels of soybean productivity to the municipality geometry. For example: are all touched pixels considered? If yes, how? What is the projection used to make compatible the raster file of FAO-GAEZ with the vector file of the municipalities? There are also small holes in some municipalities in the Northeast region that we do not know if it was treated somehow after the data were generated. To ensure that all the differences we point out are coming from different FAO-GAEZ versions and not different data handling, we compare the results across our constructed productivity variables.
We focus on the difference between soybean productivity with high and low use of inputs ($\Delta A_{j}^{soy}\coloneqq A^{soy}_{j,high}-A^{soy}_{j,low}$), which is interpreted as the productivity shock brought by the genetically modified seeds. Table 3 shows descriptive statistics of the changes in productivity. On average, version 4 measures a higher soybean productivity of 170 kg per hectare. Overall, version 4 has lower spatial heterogeneity than version 3.
@@
Table 4 shows the estimated results of Equation (9). We focus on Column (2) of Table 9 of the original article, which reports the impact of agricultural productivity on employment share in manufacturing. Column (1) in Table 4 shows the result as reported by Bustos et al. (2016), which shows that an increase of one standard deviation in potential soy productivity leads to an increase of 1.8 p.p on the share of manufacturing employment.
@@
Notes: Observations are at the municipality level. All results use a baseline sample of 4,149 municipalities. Changes in dependent variables are calculated over the years 2000 and 2010. All municipality controls are from the population census of 1991: share rural population, log of income per capita, log of population density, literacy rate. Column (1) reports the result of Bustos et al. (2016)’s Column 2 in Table 9, which uses soybean suitability from FAO-GAEZ version 3. To make results more comparable, Columns (2) and (3) report our results using FAO GAEZ data versions 3 and 4, respectively. Columns (4) and (5) show the result when $\Delta A_{jt}^{soy}$ is normalized to have variance equal to one. Column (6) presents our partial identification approach. Standard errors are robust to heteroskedasticity. Confidence intervals in Column (6) are built at the 10% and 1% levels.
@@
We can now discuss the results of our proposed econometric tool. To make the comparison easier, Columns (4) and (5) show the result of estimating Expression (9) with FAO-GAEZ versions 3 and 4 after normalizing the main explanatory variable. In Columns (4) and (5), the interpretation of the estimated results is that an increase of one standard deviation in $\Delta A_{jt}^{soy}$ causes an increase of 1 p.p and 1.4 p.p in the share of manufacturing jobs.
@@
We also report in Column (6) the confidence intervals at the 10% and 1% levels. Even at the 10%-confidence level, we see that the lower bound is very close to the point estimates associated with OLS regressions using FAO GAEZ versions 3 and 4 (Columns (4) and (5)). Additionally, when ignoring measurement error, Bustos et al. (2016)’s result and our results in Columns (2)-(5) are all significant at the 1% level. Interestingly, when we account for measurement error, our partial identification approach still rejects the null of no effect at the 1%-confidence level. This results suggests that Bustos et al. (2016)’s result is robust to measurement error.
@@
Acharya et al. (2016) study the impact of slavery on contemporary political attitudes in the South of the United States. In particular the authors estimate the following equation:
$y_{j}=\beta\cdot S_{j}+\sum_{l=1}^{L}\alpha_{l}\cdot W_{l,j}+\epsilon_{j},$
where $y_{j}$ is the proportion of republicans in the white population in county $j$ pooled from Cooperative Congressional Election Study for the years 2006, 2008, 2009, 2010 and 2011, and $S_{j}$ is the proportion of slaves in county $j$ in 1860. The authors point out at least two potential problems with estimating the above expression via OLS: measurement error in the explanatory variable and omitted variable bias. To overcome these concerns the authors propose to use the suitability of cotton from FAO-GAEZ data version 3 as an instrument for slavery $(S_{j})$. To fit this approach within our framework, we focus on the reduced-form specification given by Equation (10):
$y_{j}=\beta\cdot A_{j}^{cotton}+\sum_{l=1}^{L}\alpha_{l}\cdot W_{l,j}+\epsilon_{j}$ (10)
The authors use an average of the FAO-GAEZ version 3 with high and intermediate inputs. As FAO-GAEZ version 4 has no intermediate input specification, we reproduce their paper using the high input specification.
@@ Conclusion
In this paper, we develop a novel partial identification econometric method to identify the effect of a mismeasured variable when two proxies are available. Importantly, our estimator relies on assumptions no stronger than those commonly applied in linear models that ignore measurement error. Moreover, our partial identification approach provides bounds that (i) incorporate all available information from the first two moments of the data distribution and (ii) compare favorably against alternative approaches, such as a simple reduced-form estimand and the method proposed by Lubotsky and Wittenberg (2006).
@@ A.1.1 Data Restrictions
$=0$ because $\text{because }X_{res}^{*}\text{ and }\epsilon_{res}\text{ are residuals.}$.
@@
$=\mathbb{E}\left[U_{k}\right]$ because $\text{because }X_{res}^{*}\text{ is a residual.}$.
@@ A.1.3 No Valid Data-Generating Process if $LB>UB$
because $\hskip 20.0pt\text{because }\tilde{\beta}\leq UB$
@@
$\hskip 40.0pt+\left[1+Cov\left(Z_{1,res},Z_{2,res}\right)-\left[\dfrac{Cov\left(Z_{1,res},Y_{res}\right)+Cov\left(Z_{2,res},Y_{res}\right)}{\tilde{\beta}}\right]\right]$
@@ Appendix B Simple Examples
where $\hskip 20.0pt\text{where }\dfrac{1+Cov\left(X_{res}^{*},U_{1}\right)}{1+Var\left(U_{1}\right)+2\cdot Cov\left(X_{res}^{*},U_{1}\right)}>0$
because $\hskip 20.0pt\text{because }Cov\left(X_{res}^{*},U_{1}\right)>-\dfrac{1+Var\left(U_{1}\right)}{2}>-1$
@@
because $\hskip 20.0pt\text{because }Cov\left(X_{res}^{*},U_{1}\right)<-Var\left(U_{1}\right)$
@@ C.1.3 No Valid Data-Generating Process if $LB_{C}>UB$
because $\hskip 20.0pt\text{because }\tilde{\beta}\leq UB$
@@
$\hskip 40.0pt+\left[1+Cov\left(Z_{1,res},Z_{2,res}\right)-\left[\dfrac{Cov\left(Z_{1,res},Y_{res}\right)+Cov\left(Z_{2,res},Y_{res}\right)}{\tilde{\beta}}\right]\right]$
@@ Appendix D No Prior Knowledge of the Target Parameter’s Sign
The reduced-form coefficients identify the sign of the true target parameter, i.e., $sign\left\{Cov\left(Z_{1},Y_{res}\right)\right\}=sign\left\{Cov\left(Z_{2},Y_{res}\right)\right\}=sign\left\{\beta\right\}$. [^15] The $sign$ function equals one if its argument is a positive number, zero if its argument is zero, and minus one if its argument is a strictly negative number.
@@ Appendix E Estimation and Inference: Detailed Explanation
      Calculate $Z_{r}^{*}\left(k\right)\coloneqq\frac{\widehat{g}_{k}^{\prime}\cdot Z_{r}}{\left|\left|\widehat{g}_{k}\right|\right|}$ for $r\in\left\{1,\ldots,R\right\}$.
@@
      Compute $\widehat{K}\coloneqq\left\{k\in\left\{1,2,3\right\}\colon\widehat{l}_{k}\geq\left(\max\limits_{\tilde{k}\in\left\{1,2,3\right\}}\left\{\widehat{l}_{\tilde{k}}-\kappa\cdot\widehat{s}_{\tilde{k}}\right\}-2\cdot\kappa\cdot\widehat{s}_{k}\right)\right\}$.
@@ Appendix F Applications Using Other Datasets
First, many recent studies (Olken, 2009; Adena et al., 2015) focus on the impact of media access on economic behavior. To do so, they frequently measure media access by combining an irregular terrain model and the location of TV or radio transmitters. Since the most commonly used irregular terrain model is in its third version (Hufford, 1995), this setting is very similar to the FAO-GAEZ productivity measures discussed in Sections 2.1 and 5.
@@
Second, many articles (Post and Hanewald, 2013; Gan et al., 2015; Bissonnette et al., 2017) discuss the impact of subjective beliefs about life expectancy on economic behavior (e.g., savings, investment, bequests). To measure this type of belief, many surveys ask individuals about their subjective probability of attaining various ages. However, individual responses frequently present measurement issues, such as rounding and focal answers. To address these problems, many authors adopt different imputation methods that approximate the true subjective probability of attaining a certain age. By treating the results of each imputation model as a different proxy, our proposed method may be used in this empirical setting.
@@
Third, multiple studies (Hsieh and Klenow, 2009, 2010) are interested in the impact of total factor productivity on economic outcomes, such as plant exit and income levels. To measure total factor productivity, different authors adopt different models (Olley and Pakes, 1996; Hsieh and Klenow, 2009; Ackerberg et al., 2015). By treating the measures produced by each model as a different proxy for the true latent total factor productivity, our partial identification strategy may be used in this empirical context.
@@
Fourth, economic historians (Sala-i-Martin, 1996; Hartley, 2021) frequently focus on understanding GDP growth and fluctuations in the long run. To do so, they need measures of income levels over long periods of time, including moments when national accounts were not collected in a systematic way. To access this type of variable, most authors use the Maddison Project Data (Bolt and van Zanden, 2025), which is in its fifth version. Each version varies either because new countries were included in the sample or, more importantly to us, new information was collected about existing countries and their historical income data were updated. Since different versions of these variables may have different measurement error terms when compared to the true income level of each country, this setting is very similar to the FAO-GAEZ productivity measures discussed in Sections 2.1 and 5.
@@
Fifth, Rodrik and Wacziarg (2005) analyze the impact of democratic transitions on economic outcomes. To measure their treatment variable, they use the Polity IV index. This expert-coded measure of democracy levels is in its fourth version and is based on experts’ interpretation of political regimes in each country and each year. Since different experts might interpret the same political regime differently, there is measurement error in each version of this index. Although this setting is similar to the FAO-GAEZ productivity measures (Sections 2.1 and 5), we believe that Assumption 5 may not be appropriate in this context. Since political experts might overreact to reanalyses of historical phenomena, it may be the case that the measurement error terms in different versions of the Polity Index are negatively correlated. In this case, the bounds proposed in Appendix C represent a more appropriate approach than the bounds in the main text.
@@
Sixth, multiple articles (Alfaro et al., 2008; Rajkumar and Swaroop, 2008; Aksoy et al., 2025) are interested in the impact of a country’s governance quality on economic outcomes, such as health quality, political trust and international capital flows. To measure governance quality, Alfaro et al. (2008) and Rajkumar and Swaroop (2008) use a variety of measures from the Political Risk Services Group, while Aksoy et al. (2025) uses the V-DEM index and one measure from the Worldwide Governance Indicators. By treating these different measures as different proxies for the true unobservable governance quality, our partial identification strategy may be applied in this empirical context.
@@ Appendix H Assumptions on Model Primitives
Our second assumption on the primitives of our model imposes a restriction on the variance of true unobservable measure.
Assumption H.2
$Var\left(X^{*}\right)=1+Var\left(W\cdot\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot X^{*}\right]\right)$.
Note that
$Var\left(X^{*}\right)$ $=Var\left(X_{res}^{*}+W\cdot\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot X^{*}\right]\right)$
$=Var\left(X_{res}^{*}\right)+Var\left(W\cdot\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot X^{*}\right]\right)$
$\hskip 20.0pt+2\cdot Cov\left(X_{res}^{*},W\right)\cdot\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot X^{*}\right]$
$=Var\left(X_{res}^{*}\right)+Var\left(W\cdot\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot X^{*}\right]\right).$
Hence, Assumption H.2 implies Assumption 2
@@
where $\hskip 20.0pt\text{where }\delta_{k}\coloneqq\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot Z_{k}\right]$
and $\hskip 20.0pt\text{and }\delta_{X}\coloneqq\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot X^{*}\right]$
@@
Hence, Assumptions H.1 and H.3 implies Assumption 4.

### v4

@@ Introduction
$\Delta y_{it}=\alpha_{t}+\beta\cdot\Delta A_{it}^{soy}+\sum_{j=1}^{J}\alpha_{j}\cdot w_{i,j}+\epsilon_{jt},$ (9)
where $y_{it}$ is the employment share of the manufacturing sector in municipality $i$ and year $t$, and $\Delta A_{it}^{soy}$ is the soybean productivity shock, which is the variable of interest $X^{*}$. Bustos et al. (2016) exploits an arguably exogenous shock in soybean productivity: the introduction of genetically modified seeds in 2003, i.e., they use this increase in productivity as an exogenous change in soybean productivity.
@@
We begin by discussing the differences in soybean productivity in versions 3 and 4. We are able to reproduce the numbers of productivity using Bustos et al. (2016)’s replication package. The correlation between the change in productivity variable in the replication package and the one generated by us is of 0.99.151515There are many reasons for differences in the computed productivity. In the replication package, there is no mention of how the variables were transformed from raster and vector formats to the tabular format the authors use. We use the geometry of the minimum comparable areas that is present in the replication package in order to compute the average productivity within each municipality. The replication package does not have the raw soybean productivity data, but we know of no different vintages of the FAO-GAEZ version 3 and, therefore, we have no reason to believe that the data is different. The differences then are likely to occur from different ways of assigning pixels of soybean productivity to the municipality geometry. For example: are all touched pixels considered? If yes, how? What is the projection used to make compatible the raster file of FAO-GAEZ with the vector file of the municipalities? There are also small holes in some municipalities in the Northeast region that we do not know if it was treated somehow after the data were generated. To ensure that all the differences we point out are coming from different FAO-GAEZ versions and not different data handling, we compare the results across our constructed productivity variables.
We focus on the difference between soybean productivity with high and low use of inputs ($\Delta A_{i}^{soy}\coloneqq A^{soy}_{i,high}-A^{soy}_{i,low}$), which is interpreted as the productivity shock brought by the genetically modified seeds. In our setting, the variable $\Delta A_{i}^{soy}$ is the variable of interest $X^{*}$. Table 3 shows descriptive statistics of the changes in productivity. On average, version 4 measures a higher soybean productivity of 170 kg per hectare. Overall, version 4 has lower spatial heterogeneity than version 3.
@@
Table 4 shows the estimated results of Equation (9). We focus on Column (2) of Table 9 of the original article, which reports the impact of agricultural productivity on employment share in manufacturing. Column (1) in Table 4 shows the result as reported by Bustos et al. (2016), which shows that an increase of one standard deviation in potential soy productivity leads to an increase of 1.7 p.p $0.83\times 0.021$] on the share of manufacturing employment.
@@
Notes: Observations are at the municipality level. All results use a baseline sample of 4,149 municipalities. Changes in dependent variables are calculated over the years 2000 and 2010. All municipality controls are from the population census of 1991: share rural population, log of income per capita, log of population density, literacy rate. Column (1) reports the result of Bustos et al. (2016)’s Column 2 in Table 9, which uses soybean suitability from FAO-GAEZ version 3. To make results more comparable, Columns (2) and (3) report our results using FAO-GAEZ data versions 3 and 4, respectively. Columns (4) and (5) show the result when $\Delta A_{it}^{soy}$ is normalized to have variance equal to one. Column (6) presents our partial identification approach. Standard errors are robust to heteroskedasticity. Confidence intervals in Column (6) are built at the 10% and 1% levels.
@@
We can now discuss the results of our proposed econometric tool. To make the comparison easier, Columns (4) and (5) show the result of estimating Expression (9) with FAO-GAEZ versions 3 and 4 after normalizing the main explanatory variable. In Columns (4) and (5), the interpretation of the estimated results is that an increase of one standard deviation in $\Delta A_{it}^{soy}$ causes an increase of 1 p.p and 1.4 p.p in the share of manufacturing jobs.
@@
We also report in Column (6) the confidence intervals at the 10% and 1% levels. Even at the 10%-confidence level, we see that the lower bound is very close to the point estimates associated with OLS regressions using FAO-GAEZ versions 3 and 4 (Columns (4) and (5)). Additionally, when ignoring measurement error, Bustos et al. (2016)’s result and our results in Columns (2)-(5) are all significant at the 1% level. Interestingly, when we account for measurement error, our partial identification approach still rejects the null of no effect at the 1%-confidence level. This results suggests that Bustos et al. (2016)’s result is robust to measurement error.
@@
Acharya et al. (2016) study the impact of slavery on contemporary political attitudes in the South of the United States. In particular, the authors estimate the following equation:
$y_{i}=\beta\cdot S_{i}+\sum_{j=1}^{J}\alpha_{j}\cdot w_{i,j}+\epsilon_{j},$
where $y_{i}$ is the proportion of republicans in the white population in county $i$ pooled from Cooperative Congressional Election Study for the years 2006, 2008, 2009, 2010 and 2011, and $S_{i}$ is the proportion of slaves in county $i$ in 1860. The authors point out at least two potential problems with estimating the above expression via OLS: measurement error in the explanatory variable and omitted variable bias. To overcome these concerns the authors propose to use the suitability of cotton from FAO-GAEZ data version 3 as an instrument for slavery $(S_{i})$. To fit this approach within our framework, we focus on the reduced-form specification given by Equation (10):
$y_{i}=\beta\cdot A_{i}^{cotton}+\sum_{j=1}^{J}\alpha_{j}\cdot W_{i,j}+\epsilon_{i}$ (10)
The authors use an average of the FAO-GAEZ version 3 with high and intermediate inputs. As FAO-GAEZ version 4 has no intermediate input specification, we reproduce their paper using the high input specification. In our setting, the variable $A_{i}^{cotton}$ is the variable of interest $X^{*}$.
@@ Conclusion
In this paper, we develop a novel partial identification analytical method to identify the effect of a mismeasured variable when two proxies are available. Importantly, our estimator relies on assumptions no stronger than those commonly applied in linear models that ignore measurement error. Moreover, our partial identification approach provides bounds that (i) incorporate all available information from the first two moments of the data distribution and (ii) compare favorably against alternative approaches, such as a simple reduced-form estimand and the method proposed by Lubotsky and Wittenberg (2006).
@@ A.1.1 Data Restrictions
$=0$ $\text{because }X_{res}^{*}\text{ and }\epsilon_{res}\text{ are residuals.}$
@@
$=\mathbb{E}\left[U_{k}\right]$ $\text{because }X_{res}^{*}\text{ is a residual.}$
@@ A.1.3 No Valid Data-Generating Process if $LB>UB$
$\hskip 20.0pt\text{because }\tilde{\beta}\leq UB$
@@
[…]
@@ Appendix B Simple Examples
$\hskip 20.0pt\text{where }\dfrac{1+Cov\left(X_{res}^{*},U_{1}\right)}{1+Var\left(U_{1}\right)+2\cdot Cov\left(X_{res}^{*},U_{1}\right)}>0$
$\hskip 20.0pt\text{because }Cov\left(X_{res}^{*},U_{1}\right)>-\dfrac{1+Var\left(U_{1}\right)}{2}>-1$
@@
$\hskip 20.0pt\text{because }Cov\left(X_{res}^{*},U_{1}\right)<-Var\left(U_{1}\right)$
@@ C.1.3 No Valid Data-Generating Process if $LB_{C}>UB$
$\hskip 20.0pt\text{because }\tilde{\beta}\leq UB$
@@
[…]
@@ Appendix D No Prior Knowledge of the Target Parameter’s Sign
The reduced-form coefficients identify the sign of the true target parameter, i.e., $sign\left\{Cov\left(Z_{1},Y_{res}\right)\right\}=sign\left\{Cov\left(Z_{2},Y_{res}\right)\right\}=sign\left\{\beta\right\}$. [^17] The signsign function equals one if its argument is a strictly positive number, zero if its argument is zero, and minus one if its argument is a strictly negative number.
@@ Appendix E Estimation and Inference: Detailed Explanation
[…]
@@
[…]
@@ Appendix F Applications Using Other Datasets
First, many recent studies (Olken, 2009; Adena et al., 2015) focus on the impact of media access on economic behavior. To do so, they frequently measure media access by combining an irregular terrain model and the location of TV or radio transmitters. Since the most commonly used irregular terrain model is in its third version (Hufford, 1995), this setting is very similar to the FAO-GAEZ productivity measures discussed in Sections 2.1 and 5. For example, we can use our proposed tool to analyze a reduced-form version of Panel B in Table 11 by Olken (2009) and reanalyze Equation (1) by Adena et al. (2015).
@@
Second, many articles (Post and Hanewald, 2013; Gan et al., 2015; Bissonnette et al., 2017) discuss the impact of subjective beliefs about life expectancy on economic behavior (e.g., savings, investment, bequests). To measure this type of belief, many surveys ask individuals about their subjective probability of attaining various ages. However, individual responses frequently present measurement issues, such as rounding and focal answers. To address these problems, many authors adopt different imputation methods that approximate the true subjective probability of attaining a certain age. By treating the results of each imputation model as a different proxy, our proposed method may be used in this empirical setting. For example, we can reanalyze Equation (4-3) by Post and Hanewald (2013).
@@
Third, multiple studies (Hsieh and Klenow, 2009, 2010) are interested in the impact of total factor productivity on economic outcomes, such as plant exit and income levels. To measure total factor productivity, different authors adopt different models (Olley and Pakes, 1996; Hsieh and Klenow, 2009; Ackerberg et al., 2015). By treating the measures produced by each model as a different proxy for the true latent total factor productivity, our partial identification strategy may be used in this empirical context. For example, we can reanalyze Table VIII by Hsieh and Klenow (2009).
@@
Fourth, economic historians (Sala-i-Martin, 1996; Hartley, 2021) frequently focus on understanding GDP growth and fluctuations in the long run. To do so, they need measures of income levels over long periods of time, including moments when national accounts were not collected in a systematic way. To access this type of variable, most authors use the Maddison Project Data (Bolt and van Zanden, 2025), which is in its fifth version. Each version varies either because new countries were included in the sample or, more importantly to us, new information was collected about existing countries and their historical income data were updated. Since different versions of these variables may have different measurement error terms when compared to the true income level of each country, this setting is very similar to the FAO-GAEZ productivity measures discussed in Sections 2.1 and 5. For example, we can use our proposed tool to reanalyze Equations (1), (3) and (5) by Sala-i-Martin (1996) and Equations (1) and (2) by Hartley (2021).
@@
Fifth, Rodrik and Wacziarg (2005, Tables 1 and 2) analyze the impact of democratic transitions on economic outcomes. To measure their treatment variable, they use the Polity IV index. This expert-coded measure of democracy levels is in its fourth version and is based on experts’ interpretation of political regimes in each country and each year. Since different experts might interpret the same political regime differently, there is measurement error in each version of this index. Although this setting is similar to the FAO-GAEZ productivity measures (Sections 2.1 and 5), we believe that Assumption 5 may not be appropriate in this context. Since political experts might overreact to reanalyses of historical phenomena, it may be the case that the measurement error terms in different versions of the Polity Index are negatively correlated. In this case, the bounds proposed in Appendix C represent a more appropriate approach than the bounds in the main text.
@@
Sixth, multiple articles (Alfaro et al., 2008; Rajkumar and Swaroop, 2008; Aksoy et al., 2025) are interested in the impact of a country’s governance quality on economic outcomes, such as health quality, education quality, political trust and international capital flows. To measure governance quality, Alfaro et al. (2008) and Rajkumar and Swaroop (2008) use a variety of measures from the Political Risk Services Group, while Aksoy et al. (2025) uses the V-DEM index and one measure from the Worldwide Governance Indicators. By treating these different measures as different proxies for the true unobservable governance quality, our partial identification strategy may be applied in this empirical context. For example, we can reanalyze Tables 5, 9 and 10 (Panel C) by Alfaro et al. (2008), Equations (1) and (2) by Rajkumar and Swaroop (2008), and Equation (2) by Aksoy et al. (2025).
@@ Appendix H Assumptions on Model Primitives
Our second assumption on the primitives of our model is Assumption 2 directly because we see it as modifying the target parameter, as explained in the main text.
@@
$\hskip 20.0pt\text{where }\delta_{k}\coloneqq\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot Z_{k}\right]$
$\hskip 20.0pt\text{and }\delta_{X}\coloneqq\left(\mathbb{E}\left[W^{T}\cdot W\right]\right)^{-1}\cdot\mathbb{E}\left[W^{T}\cdot X^{*}\right]$
@@
Hence, Assumptions H.1 and H.2 implies Assumption 4.
@@
Appendix I Sources of Measurement Error in FAO-GAEZ data
In this appendix, we discuss potential sources of measurement error in the FAO-GAEZ data. Our main goal is arguing the plausibility of Assumption 5.
Consider the following model describing how FAO-GAEZ version $v$ proxies for agricultural productivity of crop $c$
$Z_{v}=f_{v}(M+\eta_{v},\theta+\zeta_{v})$ (I.1)
where $f_{v}$ is a known function, $M$ denotes a vector of predictors (e.g., true precipitation and true temperature), and $\theta$ is a vector of parameters. Importantly, $M$ and $\theta$ are permanent objects in the sense that they do not change across versions. Since these permanent objects are not perfectly observed, we include two error terms that vary across versions: $\zeta_{v}$ captures calibration error and $\eta_{v}$ captures measurement error in the predictors.
Thus, the measurement error in version $v$ can be decomposed as:
$\begin{split}U_{v}&=f_{v}(M+\eta_{v},\theta+\zeta_{v})-X^{*}\\ &=\underbrace{f_{v}(M+\eta_{v},\theta+\zeta_{v})-f_{v}(M,\theta)}_{\text{imperfect data term}}+\underbrace{f_{v}(M,\theta)-X^{*}}_{\text{functional form misspecification}}\end{split}$ (I.2)
Intuitively, the first term $\left(f_{v}(M+\eta_{v},\theta+\zeta_{v})-f_{v}(M,\theta)\right)$ captures measurement error arising from imprecise input data. This term will be positively correlated across versions if the input data presents worse quality in developing regions over time. For example, developing regions may have a smaller number of weather stations and may rely more heavily on interpolation methods to estimate the predictors $M$, implying that $\eta_{v}$ is correlated across versions of the FAO-GAEZ project.
The second term $\left(f_{v}(M,\theta)-X^{*}\right)$ captures measurement error arising from functional form misspecification. This type of error occurs because researchers do not know the true model $f^{*}$ connecting predictors and parameters to crop productivity $X^{*}=f^{*}\left(M,\theta\right)$. This term will be positively correlated across versions if the proxy model $f_{v}$ is not updated heavily across versions of the FAO-GAEZ project. In particular, this term will be perfectly correlated across versions if $f_{v}$ is the same function for different values of $v$.
Lastly, this decomposition illustrates why Assumption 5 is plausible. A sufficient condition for $Cov(U_{1},U_{2})\geq 0$ is (i) persistence on both the imperfect data term and the misspecification term across versions and (ii) a non-negative covariance between these two terms across versions. According to the last two paragraphs, condition (i) is plausible in our setting. Furthermore, condition (ii) may arise if predictor measurement and functional-form assumptions are less accurate in developing regions. The latter is plausible in our context because scientific investment in developing regions is smaller.